OpenStudy (anonymous):

Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

OpenStudy (anonymous):

I WILL MEDAL AND FAN PLEASE HELP!

OpenStudy (anonymous):

@Hero @Ashleyisakitty

OpenStudy (anonymous):

Hi! do you know the formula for the volume of a cube? It is \(\sf V_{cube} = s^3\) where s = 10 plug it in the formula and you'll get it :)

OpenStudy (anonymous):

@Data_LG2 30 cubic units?

OpenStudy (anonymous):

nope. \(\sf s^3 = 10^3 = 10 \times 10 \times 10 = ?\)

OpenStudy (anonymous):

oh lol im dumb

OpenStudy (anonymous):

no, don't say that everybody have those times :) so what will you get?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

10 x 10 is 100. What is 100 x 10?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

What is 10^3 ?

OpenStudy (anonymous):

1000 x 3= 3000

OpenStudy (anonymous):

Made two mistaken postings and they were deleted. 3000 = 3*1000 = 3 * 10^3

OpenStudy (anonymous):

i thought u meant when u said 10^3 u meant 1000 x 3, sorry